Computer organization
roblem 4 (17 pts) ) A program runs 10 us in 10 seconds on computer A/ which has a 2 GHz clock. You are trying to help a computer designer determined that a build a computer, B, which will run this program in 6 seconds The designer has substantial increase in the clock rate is possible, but this increase will affect the rest causing computer B to require L.2 times as many clock cycles as computer A for CPU design, clock rate should you tell the designer to target for computer B? this program. What
Answer #1

Computer A : 2 GHz = 2*109 cycle per second

Program run in 10 seconds on computer A this Program requires 10*2*109 cycle

Now for computer the number of cycle is 1.2 time of original = 1.2*2*1010 =24*109 cycle

24*109 cycle it should be done in 6 second

That means in 1 second = (24*109 )/6 = 4*109 cycle

4*109 cycle per second means 4Ghz

Hence for computer B, we need 4Ghz Clock rate.

THUMBS UP IF YOU ARE SATISFIED WITH THE ANSWER OTHERWISE REPLY WITH YOUR CONCERN.